# From Discrete to Continuous Binary Best-Response Dynamics: Discrete Fluctuations Almost Surely Vanish with Population Size

### v3

@@ VII-D Proof of Lemma 1 when agents include themselves
$\begin{cases}{1}&\text{if }(x^{\mathsf{N}}\leq{\tau_{p}}\text{ and }p\leq\mathsf{p})\text{ or }(x^{\mathsf{N}}\geq\tau_{\mathsf{p}+\mathsf{p}^{\prime}+1-p}^{\prime}\text{ and }p>\mathsf{p}),\\ {2}&\text{if }(x^{\mathsf{N}}>{\tau_{p}}\ \text{ and }p\leq\mathsf{p})\text{ or }(x^{\mathsf{N}}<\tau_{\mathsf{p}+\mathsf{p}^{\prime}+1-p}^{\prime}\text{ and }p>\mathsf{p}).\end{cases}$
If Lemma 1 remains true, then the main results of the papers are also valid. Condition 1 is already satisfied. Let $\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x}^{\mathsf{N}})$ denote the expected increment per time unit $t$ of the Markov chain $\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle_{k}$ when the Markov state is at $\bm{x}^{\mathsf{N}}$, i.e., $\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x}^{\mathsf{N}})=\smash{\mathsf{N}\mathbb{E}[\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k+1}-\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\mid\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}=\bm{x}^{\mathsf{N}}]}$. The $p$th element of $\bm{\nu}^{\frac{1}{\mathsf{N}}}$, $p\in[\mathsf{p}+\mathsf{p}^{\prime}]$, reads as ${\nu}^{\frac{1}{\mathsf{N}}}_{p}({\bm{x}}^{\mathsf{N}})=\rho_{p}(2-\hat{s}(p,\bm{x}^{\mathsf{N}}))-x^{\mathsf{N}}_{p}$. Accordingly, Conditions 2 and 4 can be easily verified. Regarding Condition 3, For each given $\delta$, we need to find an $\epsilon_{0}$ which satisfies (5). In what follows, we show that (5) is satisfied for any arbitrarily chosen $\epsilon_{0}$. We consider a selection from (7) denoted by $\dot{\bm{x}}=\bm{\nu}(\bm{x})$ where $\bm{\nu}$ is a $(\mathsf{p}+\mathsf{p}^{\prime})$-dimensional function and its $p$th element ${\nu}_{p}$ equals $\rho_{p}-x_{p}$ if $\mathtt{A}$ is the preferred strategy of subpopulation $p$ and equals $-x_{p}$ if the preferred strategy of subpopulation $p$ is $\mathtt{B}$. This results in ${\nu}_{p}(\bm{x})=\rho_{p}(2-\hat{s}(p,\bm{x}))-x_{p}$. The expected increment per unit of time, $\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x})$, is the same as the function $\bm{\nu}(\bm{x})$. As a result, equation (5) and, in turn, Condition 3 are satisfied by selecting $\bm{y}$ equal to $\bm{x}$, regardless of the value of $\epsilon_{0}$. This implies that Lemma 1 holds.
@@ VII-E Proof of 2
It holds that $x=\Sigma_{k=1}^{\mathsf{p}+\mathsf{p}^{\prime}}x_{k}$ and $0\leq x\leq 1$. Considering 1, the thresholds of the subpopulations are distinct, building $\mathsf{p}+\mathsf{p}^{\prime}+1$ disjoint open intervals in the unit interval, where their limit points are either 00, $1$, or the thresholds. Denote by $\mathcal{T}$ the union of the set $\{0,1\}$ with the set of all subpopulations’ thresholds. When $x\notin\mathcal{T}$, there exist $i\in\{0,1,\ldots,\mathsf{p}+1\}$ and $j\in\{0,1,\ldots,\mathsf{p}^{\prime}+1\}$ where $x\in(\tau^{\prime}_{j},\tau_{i})$ such that $\tau_{i+1}<x<\tau^{\prime}_{j+1}$. In view of (7), in this case, $\bm{\mathcal{V}}(\bm{x})$ is a singleton and for all $0<l\leq i$ anticoordinating subpopulations, $\dot{x}_{l}=\rho_{l}-x_{l}$ and for the remaining anticoordinating subpopulations, i.e., $i<l\leq\mathsf{p}$, $\dot{x}_{l}=-x_{l}$. Similarly, for coordinating subpopulations, $\dot{x}_{l}=-x_{l}$ for all $\mathsf{p}<l\leq\mathsf{p}+\mathsf{p}^{\prime}-j+1$ and $\dot{x}_{l}=\rho_{l}-x_{l}$ for all $\mathsf{p}+\mathsf{p}^{\prime}-j+1<l\leq\mathsf{p}+\mathsf{p}^{\prime}$. By summing over the elements of $\dot{\bm{x}}$, we have $\dot{x}=\Sigma_{k=1}^{\mathsf{p}+\mathsf{p}^{\prime}}\dot{x}_{k}={\eta}_{i}+{\eta}^{\prime}_{j}-x$. Therefore, for $x\in(\tau^{\prime}_{j},\tau_{i})$, $\mathcal{X}(x)=\{{\eta}_{i}+{\eta}^{\prime}_{j}-x\}$ which is the same as the last case of $\mathcal{X}$ in (11). When $x\in\mathcal{T}$, three cases may happen: $x=\tau^{\prime}_{j}$, $x=\tau_{i}$, or $x\in\{0,1\}$. When $x=\tau_{i}$, based on (7), except for $\mathcal{V}_{i}$, the elements of $\bm{\mathcal{V}}(\bm{x})$ are single-valued and contain either $\rho_{l}-x_{l}$ or $-x_{l}$. More specifically, for anticoordinating subpopulations, we have $\dot{x}_{l}=\rho_{l}-x_{l}$ for $l<i$ and $\dot{x}_{l}=-x_{l}$ for $i<l\leq\mathsf{p}$. As for coordinating subpopulations, assume that the set $\{k\in[\mathsf{p}^{\prime}]|\tau^{\prime}_{k}<\tau_{i}\}$ is not empty and let $j=\max\{k\in[\mathsf{p}^{\prime}]|\tau^{\prime}_{k}<\tau_{i}\}$. Then, for $l$ satisfying $\mathsf{p}+\mathsf{p}^{\prime}-j+1\leq l\leq\mathsf{p}+\mathsf{p}^{\prime}$, $\dot{x}_{l}$ will be equal to $\rho_{l}-x_{l}$, and for the remaining subpopulations will be equal to $-x_{l}$. If $\{k\in[\mathsf{p}^{\prime}]|\tau^{\prime}_{k}<\tau_{i}\}$ is empty, then $\dot{x}_{l}$ for $l>\mathsf{p}$ will be equal to $-x_{l}$. Summing over the elements of $\dot{\bm{x}}$ except the $i^{\text{th}}$ element, yields ${\eta}_{i-1}+{\eta}^{\prime}_{j}-(x-x_{i})$. As for $l=i$, we have $\dot{x}_{i}\in[-x_{i},\rho_{i}-x_{i}]$ yielding $\dot{x}\in({\eta}_{i-1}+{\eta}^{\prime}_{j}-x,{\eta}_{i}+{\eta}^{\prime}_{j}-x)$. Therefore, $\mathcal{X}(x)=[{\eta}_{i-1}+{\eta}^{\prime}_{j}-x,{\eta}_{i}+{\eta}^{\prime}_{j}-x]$, which is the same as the first case of $\mathcal{X}$ in (11). If $x=\tau^{\prime}_{j}$, a similar reasoning results in $\mathcal{X}(x)=[{\eta}_{i}+{\eta}^{\prime}_{j-1}-x,{\eta}_{i}+{\eta}^{\prime}_{j}-x]$, which is the same as the second case of $\mathcal{X}$ in (11). Finally, when $x\in\{0,1\}$, a similar reasoning yields $\dot{x}={\eta}_{\mathsf{p}}-x$ if $x=0$ and $\dot{x}={\eta}^{\prime}_{\mathsf{p}^{\prime}}-x$ if $x=1$. These two are equivalent to the first two cases of $\mathcal{X}$ in (11) as $\tau_{\mathsf{p}+1}=\tau_{0}^{\prime}=0$ and $\tau^{\prime}_{\mathsf{p}^{\prime}+1}=\tau_{0}=1$. ∎
@@ VII-F Proof of Lemma 2
The proofs are first done for the abstract dynamics. We first prove parts 2, 3, and 4. Part 2) (sufficiency) Should (12) hold, it would follow that $c^{ij}\in(\max\{\tau^{\prime}_{j},\tau_{i+1}\},\min\{\tau^{\prime}_{j+1},\tau_{i}\})$ and based on (11), $0\in\mathcal{X}(c^{ij})$, implying that $c^{ij}$ is an equilibrium of (11). (Necessity) Should $c^{ij}$ be an equilibrium of the abstract dynamics, then $0\in\mathcal{X}(c^{ij})$. One of the three cases of $\mathcal{X}$ in (11) must be active for $x=c^{ij}$. The first two cases of $\mathcal{X}$ cannot be active because of 1. So the third one must hold, implying that $c^{ij}\in(\max\{\tau^{\prime}_{j},\tau_{i+1}\},\min\{\tau^{\prime}_{j+1},\tau_{i}\})$, which is equivalent to (12). Part 3) and Part 4). Following Part 2, the sufficiency and necessity can be concluded. Part 1) In view of 1, equilibrium points of clean-cut, anticoordinator-driven, and coordinator-driven types do not match for (11). Hence, an abstract state can at most be only one of these three types. We prove by contradiction that an equilibrium point of (11) should be either of these three types. Suppose $x^{*}$ is an equilibrium point for (11) but is neither clean-cut, anticoordinator-driven, nor coordinator-driven. At $x=x^{*}$ one of the cases of $\mathcal{X}$ in (11) should be active. If the first case is active, we have $x^{*}=\tau_{i}$ and as it is an equilibrium point $0\in\mathcal{X}(x^{*})$. This results in ${\eta}_{i-1}+{\eta}^{\prime}_{j}<\tau_{i}<{\eta}_{i}+{\eta}^{\prime}_{j}$, which is equivalent to (13), but we assumed that $x^{*}$ is not an anticoordinator-driven equilibrium point. This is a contradiction and as a result the first case cannot be active. Similarly, for the remaining two cases, we reach a contradiction, and hence $x^{*}$ should be either clean-cut, anticoordinator-driven, or coordinator-driven. Now, we prove the lemma for the equilibrium points of the continuous-time population dynamics (7). The sufficiency and necessity of Part 2) can be easily shown. Part 3) (Sufficiency) Assume that (13) holds. By plugging $\mathbf{a}^{ij}$ into (7), we see that $\mathbf{0}\in\bm{\mathcal{V}}(\mathbf{a}^{ij})$ and hence $\mathbf{a}^{ij}$ is an equilibrium point for (7). (Necessity) Should $\mathbf{a}^{ij}$ be an equilibrium point for (7), $\mathbf{0}\in\bm{\mathcal{V}}(\mathbf{a}^{ij})$ and consequently ${0}\in\mathcal{V}_{i}(\mathbf{a}^{ij})$. This results in $0\in[-{a}^{ij}_{i},\rho_{i}-{a}^{ij}_{i}]$ and by plugging ${a}^{ij}_{i}=\tau_{i}-({\eta}_{i-1}+{\eta}^{\prime}_{j})$ into $[-{a}^{ij}_{i},\rho_{i}-{a}^{ij}_{i}]$, we have $0\in[-\tau_{i}+{\eta}_{i-1}+{\eta}^{\prime}_{j},\rho_{i}-\tau_{i}+{\eta}_{i-1}+{\eta}^{\prime}_{j}]$. This results in ${\eta}_{i-1}+{\eta}^{\prime}_{j}<\tau_{i}<+{\eta}_{i}+{\eta}^{\prime}_{j}$ which is equivalent to (13). Part 4) The sufficiency and necessity can be shown similar to Part 3. Part 1) We prove by contradiction that an equilibrium point of (7) should be either of these three types. Suppose $\bm{x}^{*}$ is an equilibrium point for (7) but is neither clean-cut, anticoordinator-driven, nor coordinator-driven. Now, based on the value of $x^{*}$, two cases may happen. Case 1) $x^{*}$ equals a threshold. Suppose that $x^{*}$ equals the threshold of one anticoordinating subpopulation, say $i$. From $\mathbf{0}\in\bm{\mathcal{V}}(\bm{x}^{*})$ we have $0\in\mathcal{V}_{l}(\bm{x}^{*})$ for $l\in[\mathsf{p}+\mathsf{p}^{\prime}]$, where for $l\neq i$, $\mathcal{V}_{l}(\bm{x}^{*})$ is a singleton and therefore to have $0\in\mathcal{V}_{l}(\bm{x}^{*})$, $x^{*}_{l}$ must equal $\rho_{l}$ for $l<i$ or $\mathsf{p}+\mathsf{p}^{\prime}-j+1\leq l\leq\mathsf{p}+\mathsf{p}^{\prime}$, and $x^{*}_{l}$ must equal 00, otherwise, where $j=\max\{k\in[\mathsf{p}^{\prime}]|\tau^{\prime}_{k}<\tau_{i}\}$ and $j=0$ if the set is empty. Therefore, $x^{*}-x^{*}_{i}={\eta}_{i-1}+{\eta}^{\prime}_{j}$. On the other hand, $x^{*}=\tau_{i}$ resulting in $x^{*}_{i}=\tau_{i}-{\eta}_{j}^{\prime}-{\eta}_{i-1}$ which is in turn equal to $a^{ij}_{i}$. This results in that $\bm{x}^{*}$ must be equivalent to $\mathbf{a}^{ij}$ which is a contradiction. Similarly, the case where $x^{*}$ equals the threshold of one coordinating subpopulation results in a contradiction. Therefore, $x^{*}$ cannot be equal to a threshold. Case 2) $x^{*}$ is not equal to any thresholds. In this case, $\bm{\mathcal{V}}(\bm{x}^{*})$ is a singleton and consequently, $\dot{\bm{x}}=\mathbf{0}$ which results in $\dot{x}={0}$. We showed that the equilibrium points of (11) should be either clean-cut, anticoordinator-driven, or coordinator-driven which is a contradiction. In view of Case 1 and Case 2, the assumption that an equilibrium point of (7) is neither clean-cut, anticoordinator-driven, nor coordinator-driven reaches a contradiction. ∎
@@
We prove by contradiction. Consider two consecutive clean-cut or anticoordinator-driven equilibrium points $q^{*}_{k-1}<q^{*}_{k}$. Assume on the contrary that there exist no coordinator-driven equilibrium points between $q^{*}_{k-1}$ and $q^{*}_{k}$. In view of the possible cases of $\mathcal{X}$ in (11), $\bar{f}(x)=\max\{y|y\in\mathcal{X}(x)\}$ will be negative for $x\in(q^{*}_{k-1},d^{+}_{k-1})$, where $d^{+}_{k-1}$ is lower bounded by $\min\{\bar{\tau}^{\prime}_{q_{k-1}},\bar{\tau}_{q_{k-1}}\}$, i.e., $d^{+}_{k-1}\geq\min\{\bar{\tau}^{\prime}_{q_{k-1}},\bar{\tau}_{q_{k-1}}\}$, where $\bar{\tau}^{\prime}_{q_{k-1}}=\tau^{\prime}_{j+1}$, $\bar{\tau}_{q_{k-1}}=\tau_{i}$ when $q^{*}_{k-1}$ is clean-cut, i.e., $q^{*}_{k-1}=c^{ij}$ and $\bar{\tau}^{\prime}_{q_{k-1}}={\tau}^{\prime}_{j+1}$, $\bar{\tau}_{q_{k-1}}=\tau_{i-1}$ when $q^{*}_{k-1}$ is anticoordinator-driven, i.e., $q^{*}_{k-1}=a^{ij}$. Based on the value of $\min\{\bar{\tau}^{\prime}_{q_{k-1}},\bar{\tau}_{q_{k-1}}\}$, two cases may happen: Case 1) $\min\{\bar{\tau}^{\prime}_{q_{k-1}},\bar{\tau}_{q_{k-1}}\}=\bar{\tau}_{q_{k-1}}$ and consequently for a small enough $\varepsilon>0$, $\bar{f}(\bar{\tau}_{q_{k-1}}-\varepsilon)<0$. The sign of $\bar{f}(\bar{\tau}_{q_{k-1}}+\varepsilon)$ cannot be positive as $\bar{\tau}_{q_{k-1}}$ is the threshold of an anticoordinating subpopulation and considering the cases of $\mathcal{X}$ in (11), the value of $\dot{x}$ at $x=\tau_{i}+\varepsilon$ is not greater than its value at $x=\tau_{i}-\varepsilon$ for any $i\in[\mathsf{p}]$. Therefore, the sign of $\bar{f}$ does not change around $x=\bar{\tau}_{q_{k-1}}$. Actually, no matter how many thresholds of anticoordinating subpopulations (non benchmark subpopulations) are in the interval $(q^{*}_{k-1},q^{*}_{k})$, the sign of $\bar{f}$ will not change in the neighborhood of them. So, in this case, $\bar{f}(\bar{\tau}_{q_{k-1}}+\varepsilon)$ remains negative. Case 2) $\min\{\bar{\tau}^{\prime}_{q_{k-1}},\bar{\tau}_{q_{k-1}}\}=\bar{\tau}^{\prime}_{q_{k-1}}$ and consequently for a small enough $\varepsilon>0$, $\bar{f}({\bar{\tau}^{\prime}_{q_{k-1}}}-\varepsilon)<0$ and considering the second case of $\mathcal{X}$ in (11), the sign of $\bar{f}(\cdot)$ changes in a neighborhood of a coordinating subpopulation’s threshold only if $\bar{\tau}^{\prime}_{q_{k-1}}$ satisfies (14), i.e., $\bar{\tau}^{\prime}_{q_{k-1}}$ is a coordinator-driven equilibrium point. However, we assumed that it is not the case. Hence, the thresholds of coordinating subpopulations also do not impact the value of $d^{+}_{k-1}$. Putting the results of Case 1 and Case 2 together, we conclude that $d^{+}_{k-1}=q^{*}_{k}$, i.e., the sign of $\bar{f}(x)$ remains negative for $x\in(q^{*}_{k-1},q^{*}_{k})$. On the other hand, considering (12), (13), and the cases of $\mathcal{X}$ in (11), it can be concluded that $\underline{f}(x)=\min\{y|y\in\mathcal{X}(x)\}>0$ for $x\in(d^{-}_{k},q^{*}_{k})$, where $d^{-}_{k}$ is upper bounded by $\max\{\bar{\tau}^{\prime}_{q^{\prime}_{k}},\bar{\tau}_{q_{k}}\}$, where $\bar{\tau}^{\prime}_{q^{\prime}_{k}}={\tau}^{\prime}_{n}$, $\bar{\tau}_{q_{k}}=\tau_{m+1}$ when $q^{*}_{k}$ is clean-cut, i.e., $q^{*}_{k}=c^{mn}$ and similarly $\bar{\tau}^{\prime}_{q^{\prime}_{k}}=\tau^{\prime}_{n}$, $\bar{\tau}_{q_{k}}=\tau_{m+1}$ when $q^{*}_{k}$ is anticoordinator-driven, i.e., $q^{*}_{k}=a^{mn}$ As $x$ decreases, according to the cases of $\mathcal{X}$ in (11), the change of the sign of $\underline{f}(x)$ from positive to negative can only happen if the flow passes through the coordinating subpopulation’s threshold which equals a coordinator-driven equilibrium point, i.e., satisfies (14). However, we assumed it would not be the case. Hence, we conclude that the sign of $\underline{f}(x)$ is positive for $x\in(q^{*}_{k-1},q^{*}_{k})$. This is a contradiction as we already reached the conclusion that the sign of $\bar{f}(x)$ is negative for $x\in(q^{*}_{k-1},q^{*}_{k})$. Putting these arguments together, it is concluded that between two consecutive clean-cut or anticoordinator-driven equilibrium points, ${q}^{*}_{k-1},{q}^{*}_{k}$, there exists at least one coordinator-driven equilibrium point. Similarly, it can be shown that the number of coordinator-driven equilibrium points between two consecutive clean-cut/anti-coordinator-driven equilibria does not exceed one. We denote such equilibrium by ${q}^{*}_{k-1,k}$. Following similar arguments, it is straightforward to show that the leftmost and rightmost equilibrium points are clean-cut or anticoordinator-driven. Up to now, it has been shown that between two consecutive clean-cut or anticoordinator-driven equilibrium points of the abstract dynamics, $q^{*}_{k-1}$ and $q^{*}_{k}$, there exists a coordinator-driven equilibrium point, $q^{*}_{k-1,k}$ such that $q^{*}_{k-1}<q^{*}_{k-1,k}<q^{*}_{k}$. On the other hand, based Lemma 2, there is a one-to-one map between equilibrium points of the abstract dynamics and those of the continuous-time population dynamics. Therefore, for each two ordered clean-cut or anticoordinator-driven equilibrium points of the continuous-time population dynamics $\mathbf{q}^{*}_{k-1}$ and $\mathbf{q}^{*}_{k}$, there exists a coordinator-driven equilibrium point $\mathbf{q}^{*}_{k-1,k}$ such that $\smash{q^{*}_{k-1}<q^{*}_{k-1,k}<q^{*}_{k}}$. ∎

### v4

@@ VII-D Proof of Lemma 1 when agents include themselves
[…]
If Lemma 1 remains true, then the main results of the papers are also valid. Condition 1 is already satisfied. Let $\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x}^{\mathsf{N}})$ denote the expected increment per time unit $t$ of the Markov chain $\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle_{k}$ when the Markov state is at $\bm{x}^{\mathsf{N}}$, i.e., $\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x}^{\mathsf{N}})=\smash{\mathsf{N}\mathbb{E}[\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k+1}-\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\mid\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}=\bm{x}^{\mathsf{N}}]}$. The $p$th element of $\bm{\nu}^{\frac{1}{\mathsf{N}}}$, $p\in[\mathsf{p}+\mathsf{p}^{\prime}]$, reads as ${\nu}^{\frac{1}{\mathsf{N}}}_{p}({\bm{x}}^{\mathsf{N}})=\rho_{p}(2-\hat{s}(p,\bm{x}^{\mathsf{N}}))-x^{\mathsf{N}}_{p}$. Accordingly, Conditions 2 and 4 can be easily verified. Regarding Condition 3, For each given $\delta$, we need to find an $\epsilon_{0}$ which satisfies (5). In what follows, we show that (5) is satisfied for any arbitrarily chosen $\epsilon_{0}$. We consider a selection from (7) denoted by $\dot{\bm{x}}=\bm{\nu}(\bm{x})$ where $\bm{\nu}$ is a $(\mathsf{p}+\mathsf{p}^{\prime})$-dimensional function and its $p$th element ${\nu}_{p}$ equals $\rho_{p}-x_{p}$ if $\mathtt{A}$ is the preferred strategy of subpopulation $p$ and equals $x_{p}$ if the preferred strategy of subpopulation $p$ is $\mathtt{B}$. This results in ${\nu}_{p}(\bm{x})=\rho_{p}(2-\hat{s}(p,\bm{x}))-x_{p}$. The expected increment per unit of time, $\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x})$, is the same as the function $\bm{\nu}(\bm{x})$. As a result, equation (5) and, in turn, Condition 3 are satisfied by selecting $\bm{y}$ equal to $\bm{x}$, regardless of the value of $\epsilon_{0}$. This implies that Lemma 1 holds.
@@ VII-E Proof of 2
It holds that $x=\Sigma_{k=1}^{\mathsf{p}+\mathsf{p}^{\prime}}x_{k}$ and $0\leq x\leq 1$. Considering 1, the thresholds of the subpopulations are distinct, building $\mathsf{p}+\mathsf{p}^{\prime}+1$ disjoint open intervals in the unit interval, where their limit points are either 0, 11, or the thresholds. Denote by $\mathcal{T}$ the union of the set $\{0,1\}$ with the set of all subpopulations’ thresholds. When $x\notin\mathcal{T}$, there exist $i\in\{0,1,\ldots,\mathsf{p}+1\}$ and $j\in\{0,1,\ldots,\mathsf{p}^{\prime}+1\}$ where $x\in(\tau^{\prime}_{j},\tau_{i})$ such that $\tau_{i+1}<x<\tau^{\prime}_{j+1}$. In view of (7), in this case, $\bm{\mathcal{V}}(\bm{x})$ is a singleton and for all $0<l\leq i$ anticoordinating subpopulations, $\dot{x}_{l}=\rho_{l}-x_{l}$ and for the remaining anticoordinating subpopulations, i.e., $i<l\leq\mathsf{p}$, $\dot{x}_{l}=-x_{l}$. Similarly, for coordinating subpopulations, $\dot{x}_{l}=-x_{l}$ for all $\mathsf{p}<l\leq\mathsf{p}+\mathsf{p}^{\prime}-j+1$ and $\dot{x}_{l}=\rho_{l}-x_{l}$ for all $\mathsf{p}+\mathsf{p}^{\prime}-j+1<l\leq\mathsf{p}+\mathsf{p}^{\prime}$. By summing over the elements of $\dot{\bm{x}}$, we have $\dot{x}=\Sigma_{k=1}^{\mathsf{p}+\mathsf{p}^{\prime}}\dot{x}_{k}={\eta}_{i}+{\eta}^{\prime}_{j}-x$. Therefore, for $x\in(\tau^{\prime}_{j},\tau_{i})$, $\mathcal{X}(x)=\{{\eta}_{i}+{\eta}^{\prime}_{j}-x\}$ which is the same as the last case of $\mathcal{X}$ in (11). When $x\in\mathcal{T}$, three cases may happen: $x=\tau^{\prime}_{j}$, $x=\tau_{i}$, or $x\in\{0,1\}$. When $x=\tau_{i}$, based on (7), except for $\mathcal{V}_{i}$, the elements of $\bm{\mathcal{V}}(\bm{x})$ are single-valued and contain either $\rho_{l}-x_{l}$ or $x_{l}$. More specifically, for anticoordinating subpopulations, we have $\dot{x}_{l}=\rho_{l}-x_{l}$ for $l<i$ and $\dot{x}_{l}=-x_{l}$ for $i<l\leq\mathsf{p}$. As for coordinating subpopulations, assume that the set $\{k\in[\mathsf{p}^{\prime}]|\tau^{\prime}_{k}<\tau_{i}\}$ is not empty and let $j=\max\{k\in[\mathsf{p}^{\prime}]|\tau^{\prime}_{k}<\tau_{i}\}$. Then, for $l$ satisfying $\mathsf{p}+\mathsf{p}^{\prime}-j+1\leq l\leq\mathsf{p}+\mathsf{p}^{\prime}$, $\dot{x}_{l}$ will be equal to $\rho_{l}-x_{l}$, and for the remaining subpopulations will be equal to $x_{l}$. If $\{k\in[\mathsf{p}^{\prime}]|\tau^{\prime}_{k}<\tau_{i}\}$ is empty, then $\dot{x}_{l}$ for $l>\mathsf{p}$ will be equal to $x_{l}$. Summing over the elements of $\dot{\bm{x}}$ except the $i^{\text{th}}$ element, yields ${\eta}_{i-1}+{\eta}^{\prime}_{j}-(x-x_{i})$. As for $l=i$, we have $\dot{x}_{i}\in[-x_{i},\rho_{i}-x_{i}]$ yielding $\dot{x}\in({\eta}_{i-1}+{\eta}^{\prime}_{j}-x,{\eta}_{i}+{\eta}^{\prime}_{j}-x)$. Therefore, $\mathcal{X}(x)=[{\eta}_{i-1}+{\eta}^{\prime}_{j}-x,{\eta}_{i}+{\eta}^{\prime}_{j}-x]$, which is the same as the first case of $\mathcal{X}$ in (11). If $x=\tau^{\prime}_{j}$, a similar reasoning results in $\mathcal{X}(x)=[{\eta}_{i}+{\eta}^{\prime}_{j-1}-x,{\eta}_{i}+{\eta}^{\prime}_{j}-x]$, which is the same as the second case of $\mathcal{X}$ in (11). Finally, when $x\in\{0,1\}$, a similar reasoning yields $\dot{x}={\eta}_{\mathsf{p}}-x$ if $x=0$ and $\dot{x}={\eta}^{\prime}_{\mathsf{p}^{\prime}}-x$ if $x=1$. These two are equivalent to the first two cases of $\mathcal{X}$ in (11) as $\tau_{\mathsf{p}+1}=\tau_{0}^{\prime}=0$ and $\tau^{\prime}_{\mathsf{p}^{\prime}+1}=\tau_{0}=1$. ∎
@@ VII-F Proof of Lemma 2
The proofs are first done for the abstract dynamics. We first prove parts 2, 3, and 4. Part 2) (sufficiency) Should (12) hold, it would follow that $c^{ij}\in(\max\{\tau^{\prime}_{j},\tau_{i+1}\},\min\{\tau^{\prime}_{j+1},\tau_{i}\})$ and based on (11), $0\in\mathcal{X}(c^{ij})$, implying that $c^{ij}$ is an equilibrium of (11). (Necessity) Should $c^{ij}$ be an equilibrium of the abstract dynamics, then $0\in\mathcal{X}(c^{ij})$. One of the three cases of $\mathcal{X}$ in (11) must be active for $x=c^{ij}$. The first two cases of $\mathcal{X}$ cannot be active because of 1. So the third one must hold, implying that $c^{ij}\in(\max\{\tau^{\prime}_{j},\tau_{i+1}\},\min\{\tau^{\prime}_{j+1},\tau_{i}\})$, which is equivalent to (12). Part 3) and Part 4). Following Part 2, the sufficiency and necessity can be concluded. Part 1) In view of 1, equilibrium points of clean-cut, anticoordinator-driven, and coordinator-driven types do not match for (11). Hence, an abstract state can at most be only one of these three types. We prove by contradiction that an equilibrium point of (11) should be either of these three types. Suppose $x^{*}$ is an equilibrium point for (11) but is neither clean-cut, anticoordinator-driven, nor coordinator-driven. At $x=x^{*}$ one of the cases of $\mathcal{X}$ in (11) should be active. If the first case is active, we have $x^{*}=\tau_{i}$ and as it is an equilibrium point $0\in\mathcal{X}(x^{*})$. This results in ${\eta}_{i-1}+{\eta}^{\prime}_{j}<\tau_{i}<{\eta}_{i}+{\eta}^{\prime}_{j}$, which is equivalent to (13), but we assumed that $x^{*}$ is not an anticoordinator-driven equilibrium point. This is a contradiction and as a result the first case cannot be active. Similarly, for the remaining two cases, we reach a contradiction, and hence $x^{*}$ should be either clean-cut, anticoordinator-driven, or coordinator-driven. Now, we prove the lemma for the equilibrium points of the continuous-time population dynamics (7). The sufficiency and necessity of Part 2) can be easily shown. Part 3) (Sufficiency) Assume that (13) holds. By plugging $\mathbf{a}^{ij}$ into (7), we see that $\mathbf{0}\in\bm{\mathcal{V}}(\mathbf{a}^{ij})$ and hence $\mathbf{a}^{ij}$ is an equilibrium point for (7). (Necessity) Should $\mathbf{a}^{ij}$ be an equilibrium point for (7), $\mathbf{0}\in\bm{\mathcal{V}}(\mathbf{a}^{ij})$ and consequently ${0}\in\mathcal{V}_{i}(\mathbf{a}^{ij})$. This results in $0\in[-{a}^{ij}_{i},\rho_{i}-{a}^{ij}_{i}]$ and by plugging ${a}^{ij}_{i}=\tau_{i}-({\eta}_{i-1}+{\eta}^{\prime}_{j})$ into ${a}^{ij}_{i},\rho_{i}-{a}^{ij}_{i}]$, we have $0\in[-\tau_{i}+{\eta}_{i-1}+{\eta}^{\prime}_{j},\rho_{i}-\tau_{i}+{\eta}_{i-1}+{\eta}^{\prime}_{j}]$. This results in ${\eta}_{i-1}+{\eta}^{\prime}_{j}<\tau_{i}<+{\eta}_{i}+{\eta}^{\prime}_{j}$ which is equivalent to (13). Part 4) The sufficiency and necessity can be shown similar to Part 3. Part 1) We prove by contradiction that an equilibrium point of (7) should be either of these three types. Suppose $\bm{x}^{*}$ is an equilibrium point for (7) but is neither clean-cut, anticoordinator-driven, nor coordinator-driven. Now, based on the value of $x^{*}$, two cases may happen. Case 1) $x^{*}$ equals a threshold. Suppose that $x^{*}$ equals the threshold of one anticoordinating subpopulation, say $i$. From $\mathbf{0}\in\bm{\mathcal{V}}(\bm{x}^{*})$ we have $0\in\mathcal{V}_{l}(\bm{x}^{*})$ for $l\in[\mathsf{p}+\mathsf{p}^{\prime}]$, where for $l\neq i$, $\mathcal{V}_{l}(\bm{x}^{*})$ is a singleton and therefore to have $0\in\mathcal{V}_{l}(\bm{x}^{*})$, $x^{*}_{l}$ must equal $\rho_{l}$ for $l<i$ or $\mathsf{p}+\mathsf{p}^{\prime}-j+1\leq l\leq\mathsf{p}+\mathsf{p}^{\prime}$, and $x^{*}_{l}$ must equal 0, otherwise, where $j=\max\{k\in[\mathsf{p}^{\prime}]|\tau^{\prime}_{k}<\tau_{i}\}$ and $j=0$ if the set is empty. Therefore, $x^{*}-x^{*}_{i}={\eta}_{i-1}+{\eta}^{\prime}_{j}$. On the other hand, $x^{*}=\tau_{i}$ resulting in $x^{*}_{i}=\tau_{i}-{\eta}_{j}^{\prime}-{\eta}_{i-1}$ which is in turn equal to $a^{ij}_{i}$. This results in that $\bm{x}^{*}$ must be equivalent to $\mathbf{a}^{ij}$ which is a contradiction. Similarly, the case where $x^{*}$ equals the threshold of one coordinating subpopulation results in a contradiction. Therefore, $x^{*}$ cannot be equal to a threshold. Case 2) $x^{*}$ is not equal to any thresholds. In this case, $\bm{\mathcal{V}}(\bm{x}^{*})$ is a singleton and consequently, $\dot{\bm{x}}=\mathbf{0}$ which results in $\dot{x}={0}$. We showed that the equilibrium points of (11) should be either clean-cut, anticoordinator-driven, or coordinator-driven which is a contradiction. In view of Case 1 and Case 2, the assumption that an equilibrium point of (7) is neither clean-cut, anticoordinator-driven, nor coordinator-driven reaches a contradiction. ∎
@@
[…]